7. Up to what distance can two electrons approach each other if they are moving toward each other with a relative velocity of 10^8 sm/s?

1
Expert's answer
2022-01-30T13:52:47-0500

By conservation of energy:


"2·\dfrac12mv^2=\dfrac1{4\pi\epsilon_0}\dfrac{e^2}{r},\\\space\\
r=\dfrac {e^2}{4v^2{\pi\epsilon_0m}}=2.53·10^{-10}\text{ m}."